# Weak instrumental variables due to nonlinearities in panel data: A Super Learner Control Function estimator

### v2

@@ 5 Identification: Control Function Approach
$\begin{split}E[g(\tilde{x}_{it},z_{it})|\tilde{x}_{it},z_{it},\tilde{x}_{it-1},z_{it-1}]-E[g(\tilde{x}_{it-1},z_{it-1})|\tilde{x}_{it},z_{it},\tilde{x}_{it-1},z_{it-1}]=\\ g(\tilde{x}_{it},z_{it})-g(\tilde{x}_{it-1},z_{it-1})\end{split}$ (23)
@@
$E[u_{it}|\tilde{x}_{it},z_{it},\tilde{x}_{it-1},z_{it-1}]-E[u_{it-1}|\tilde{x}_{it},z_{it},\tilde{x}_{it-1},z_{it-1}]=0$ (24)
@@
A similar analysis and conclusion is obtained using a Within transformation which is equivalent to time demeaning the data. In this case, the information set $I_{t}=\{\tilde{x}_{it-j},z_{it-j},0\leq j\leq t-1\}$.
@@ 6 Super Learner Control Function estimation
$\mathbf{T}y_{it}=\mathbf{T}x_{1it}\beta_{1_{o}}+\mathbf{T}\tilde{x}_{it}^{\prime}\beta_{2_{o}}+\rho_{o}\mathbf{T}u_{it}+\mathbf{T}\omega_{it},i\in\{1,...,N\},t\in\{t_{a},...,T\}.$ (25)
@@
We partition the set $\{1,2,...,N\}$ in B subsets $S_{1}$, $S_{2}$, …, $S_{B}$. This sample splitting is equal to the one proposed by Emmenegger and Bühlmann (2023) and it was developped without knowledge of it. Alternatively, if the cluster structure in the data is known (individuals exhibit correlation within clusters but not across them, and both cluster membership and the number of clusters are identified) the data should be partitioned accordingly. Then, we estimate $\mathbf{T}g_{o}(\tilde{x}_{it},z_{it})=E[\mathbf{T}x_{1it}|I_{t}]$ using a super learner with partition $S_{b}^{c}=\{i\in S_{g},g\neq b\}$. Then, we obtain the residuals $\widehat{\mathbf{T}u}_{it}=x_{1it}-\widehat{\mathbf{T}g_{o}(\tilde{x}_{it},z_{it})}$ for partition $S_{b}=\{i\in S_{b}\}$.
Finally, the estimator of $\theta_{o}$ for partition $S_{b}$ is the solution of the following sample moment conditions:
$\frac{1}{|S_{b}|(T-t_{a})}\sum_{i\in S_{b}}\sum_{t=t_{a}}^{T}\phi(\widehat{\mathbf{T}H}_{it};\theta,\widehat{\mathbf{T}g}_{o}(\tilde{x}_{it},z_{it}))=0,$ (26)
where $\widehat{\mathbf{T}H}_{it}=[\mathbf{T}x_{1it}\quad\mathbf{T}\tilde{x}_{it}\quad\widehat{\mathbf{T}u}_{it}]^{\prime}$, and $\widehat{\mathbf{T}u}_{it}$ are the residuals obtained in the first stage for partition $S_{b}$.
This is equivalent to estimate $\theta_{o}$ by performing ordinary least squares regression of $\mathbf{T}y_{it}$ on $\mathbf{T}x_{1it}$, $\mathbf{T}\tilde{x}_{it}$ and $\widehat{\mathbf{T}u}_{it}$ or minimizing the following quadratic loss function :
$\hat{\theta}=\begin{subarray}{c}argmin\\ \theta\end{subarray}Q_{b}(\widehat{\mathbf{T}H};\theta,\widehat{\mathbf{T}g}_{o}(\tilde{x}_{it},z_{it})),$
with $Q_{b}(\widehat{\mathbf{T}H};\theta,\widehat{\mathbf{T}g}_{o}(\tilde{x}_{it},z_{it}))=\frac{1}{|S_{b}|(T-t_{a})}\sum_{i\in S_{b}}\sum_{t=t_{a}}^{T}(\mathbf{T}y_{it}-\mathbf{T}x_{1it}\beta_{1}-\mathbf{T}\tilde{x}_{it}^{\prime}\beta_{2}-\widehat{\mathbf{T}u_{it}}\rho)^{2}$, $\tilde{X}$ a $N(T-t_{a})\times(K-1)$ vector collecting all observations of $\tilde{x}_{it}$, and $\widehat{\mathbf{T}H}$ a $|S_{b}|(T-t_{a})\times(K+1)$ matrix collecting all observations of $\mathbf{T}x_{1it}$, $\mathbf{T}\tilde{x}_{it}$, and $\widehat{\mathbf{T}u}_{it}$. Importantly, ordinary least squares estimation is not the most efficient if $T>2$, or if the $\varepsilon_{it}$ is not homoskedastic neither uncorrelated across both $i$ and $t$. In this situation, a most efficient estimation is Generalized Least Squares estimation or another appropriate estimation method that accounts for this.
@@
    Estimate the conditional expectation $E[\mathbf{T}x_{1it}|I_{t}]$ using a super learner and data correspondig to $S_{b}^{c}$.
    Predict $E[\mathbf{T}x_{1it}|I_{t}]$ for $i\in S_{b}$.
    Obtain the residuals $u_{it}$ for $i\in S_{b}$ using the prediction of $E[\mathbf{T}x_{1it}|I_{t}]$ of previous step.
@@
First-difference transformation (continuation):
@@
$\Delta y_{it}=\Delta x_{1it}\beta_{1_{o}}+\Delta\tilde{x}_{it}^{\prime}\beta_{2_{o}}+\rho_{o}\Delta u_{it}+\Delta\omega_{it},i\in\{1,...,N\},t\in\{2,...,T\}.$ (27)
First, we partition the set $\{1,2,...,N\}$ in B subsets $S_{1}$, $S_{2}$, …, $S_{B}$, and we estimate $E[\Delta x_{1it}|\tilde{x}_{it},\tilde{x}_{it-1},z_{it},z_{it-1}]$ using a super learner with partition $S_{b}^{c}$. Then, we obtain the residuals $\widehat{\Delta u}_{it}=\Delta x_{1it}-\widehat{\Delta g_{o}(\tilde{x}_{it},z_{it})}$ for partition $S_{b}$. We obtain the estimator of $\theta_{o}$ is the solution of the following sample moment conditions:
$\frac{1}{|S_{b}|(T-1)}\sum_{i\in S_{b}}\sum_{t=2}^{T}\phi(\widehat{\Delta H}_{it};\theta,\widehat{\Delta g}_{o}(\tilde{x}_{it},z_{it}))=0,$ (28)
where $\widehat{\Delta H}_{it}=[\Delta x_{1it}\quad\Delta\tilde{x}_{it}\quad\widehat{\Delta u}_{it}]^{\prime}$, and $\widehat{\Delta u}_{it}$ are the residuals obtained in the first stage for partition $S_{b}$.
@@
    Estimate the conditional expectation $E[\Delta x_{1it}|\tilde{x}_{it},\tilde{x}_{it-1},z_{it},z_{it-1}]$ using a super learner and data correspondig to $I_{b}^{c}$.
    Predict $E[\Delta x_{1it}|\tilde{x}_{it},\tilde{x}_{it-1},z_{it},z_{it-1}]$ for $i\in S_{b}$.
    Obtain the residuals $u_{it}$ for $i\in I_{b}$ using the prediction of $E[\Delta x_{1it}|\tilde{x}_{it},\tilde{x}_{it-1},z_{it},z_{it-1}]$ of previous step.
@@ 7.2 The results
In this section, we present the results of the simulation experiment. In figure 2, we present the average estimated parameter of interest $\beta_{1o}$ for different values of $a$. The methods are Within OLS (WOLS), Within 2SLS (W2SLS) using polynomial of degree 1 of the covariate and the instrumental variables, Within 2SLS using polynomial of degree 5 of the covariate and the instrumental variables (W2SLS_poly), the Super Learner Control Function estimator on the First-Differenced data (FDSLCF), and the Super Learner Control Function estimator on the Within-transformed data (WSLCF). In addition, we plot the average correlation between the endogenous covariate and the instrumental variable. We also plot the p-value of the estimated control functions using the first-differenced and the within transformed data. As $a$ is larger, the nonlinearity in the reduced form equation is more important. This translates in a lower correlation between the endogenous regressor and the instrumental variables. As a result, the Within-2SLS estimators presented a decay in their performance. In contrast, the Super Learner Control Function estimators remain stable around the true value of the parameter of interest $\beta_{1o}=1$.
@@
As an illustration, we estimate the causal effect of air pollution on student performance in the United States. The data on air pollution are obtained from Deryugina et al. (2019). This dataset presents daily data on air pollutants: PM 2.5, ozone ($O_{3}$), carbon monoxide (CO), sulfur dioxide ($SO_{2}$), nitrogen dioxide ($NO_{2}$). It also contains data on wind direction, wind speed, and weather conditions at county level. We aggregate the data to yearly level by averaging across days. The data on student performance are obtained from Stanford Educational Data Archive (SEDA) Fahle et al. (2024) for county-yearly level. We merge both datasets by county-year indicator, and keep cohort 2006. Our final sample is composed of years 2009 to 2013 and 3065 counties.
We specify the following linear regression for subject district $d$, at year $t$:
@@ 9 Conclusions
This paper proposes a triangular simultaneous equation model for panel data with additive separable individual-specific fixed effects composed of a linear structural equation with a nonlinear reduced form equation to deal with the issue of weak instrumental variables due to nonlinearities in panel data settings. The parameter of interest is the structural parameter of the endogenous variable. The identification of this parameter is obtained under the assumption of available exclusion restrictions and using a control function approach. Estimating the parameter of interest is done using an estimator that we call Super Learner Control Function estimator (SLCFE). The estimation procedure is composed of two main steps and sample splitting. We estimate the control function using a super learner using sample splitting. In the following step, we use the estimated control function to control for endogeneity in the structural equation. Sample splitting is done across the individual dimension. We perform a Monte Carlo simulation to test the performance of the estimators proposed. We conclude that the estimators perform well, provided that we can learn accurately the nuisance parameter in the first stage.

### v3

@@ 5 Identification: Control Function Approach
$\begin{split}\mathbb{E}[g(\tilde{x}_{it},z_{it})|\tilde{x}_{it},z_{it},\tilde{x}_{it-1},z_{it-1}]-\mathbb{E}[g(\tilde{x}_{it-1},z_{it-1})|\tilde{x}_{it},z_{it},\tilde{x}_{it-1},z_{it-1}]=\\ g(\tilde{x}_{it},z_{it})-g(\tilde{x}_{it-1},z_{it-1})\end{split}$ (23)
@@
$\mathbb{E}[u_{it}|\tilde{x}_{it},z_{it},\tilde{x}_{it-1},z_{it-1}]-\mathbb{E}[u_{it-1}|\tilde{x}_{it},z_{it},\tilde{x}_{it-1},z_{it-1}]=0$ (24)
@@
A similar analysis and conclusion is obtained using a Within transformation which is equivalent to time demeaning the data. In this case, the information set $I_{t}=\{\tilde{x}_{it-j},z_{it-j},-T+t\leq j\leq t-1\}$.
@@ 6 Super Learner Control Function estimation
$\mathbf{T}y_{it}=\mathbf{T}x_{1it}\beta_{1_{o}}+\mathbf{T}\tilde{x}_{it}^{\prime}\beta_{2_{o}}+\rho_{o}\mathbf{T}u_{it}+\mathbf{T}\omega_{it},i\in\{1,...,N\},t\in\{t_{a},...,T_{i}\}.$ (25)
@@
We partition the set $\{1,2,...,N\}$ in B subsets $S_{1}$, $S_{2}$, …, $S_{B}$, and denote $n_{T,b}=\sum_{i\in S_{b}}T_{i}$. This sample splitting is equal to the one proposed by Emmenegger and Bühlmann (2023) and it was developed without knowledge of it. Alternatively, if the cluster structure in the data is known (individuals exhibit correlation within clusters but not across them, and both cluster membership and the number of clusters are identified) the data should be partitioned accordingly. Then, we estimate $\mathbf{T}g_{o}(\tilde{x}_{it},z_{it})=\mathbb{E}[\mathbf{T}x_{1it}|I_{t}]$ using a super learner with partition $S_{b}^{c}=\{i\in S_{g},g\neq b\}$, and we call the estimation $\widehat{\mathbf{T}g_{o}}^{S_{b}^{c}}$. Then, we obtain the residuals $\widehat{\mathbf{T}u}_{it}^{S_{b}}=\mathbf{T}x_{1it}-\widehat{\mathbf{T}g_{o}^{S_{b}^{c}}(\tilde{x}_{it},z_{it})}$ for partition $S_{b}=\{i\in S_{b}\}$.
Finally, the estimator of $\theta_{o}$ for partition $S_{b}$ is the solution of the sample moment conditions such that:
$\frac{1}{n_{T,b}}\sum_{i\in S_{b}}\phi(O_{i};\hat{\theta}_{b},\widehat{\mathbf{T}g_{o}}^{S_{b}^{c}})=0,$ (26)
where $\phi(O_{i};\hat{\theta}_{b},\widehat{\mathbf{T}g_{o}}^{S_{b}^{c}})=(\tilde{H}_{i}^{S_{b}})^{\prime}V_{i}^{-1}\mathbb{T}_{i}(y_{i}-\tilde{H_{i}}^{S_{b}}\hat{\theta}_{b})$, $\tilde{H}_{i}^{S_{b}}=[\mathbb{T}_{i}x_{1i}\quad\mathbb{T}_{i}\tilde{X}_{i}\quad\widehat{\mathbf{T}u}_{i}^{S_{b}}]$, $\widehat{\mathbf{T}u}_{i}^{S_{b}}$ is the vector stacking the residuals obtained in the first stage for partition $S_{b}$, $\tilde{X}_{i}$ a $T_{i}\times(K-1)$ vector collecting all time observations of $\tilde{X}_{it}$.
This is equivalent to estimate $\theta_{o}$ by performing generalized least squares regression of $\mathbf{T}y_{it}$ on $\mathbf{T}x_{1it}$, $\mathbf{T}\tilde{x}_{it}$ and $\widehat{\mathbf{T}u}_{it}^{S_{b}}$ or minimizing the following quadratic loss function :
$\hat{\theta}_{b}=\begin{subarray}{c}argmin\\ \theta\end{subarray}Q_{b}(O_{i};\theta,\widehat{\mathbf{T}g_{o}}^{S_{b}^{c}}),$
with $Q_{b}(O_{i};\theta,\widehat{\mathbf{T}g_{o}}^{S_{b}^{c}})=\frac{1}{n_{T,b}}\sum_{i\in S_{b}}(\mathbb{T}_{i}y_{i}-\mathbb{T}_{i}x_{1i}\beta_{1}-\mathbb{T}_{i}\tilde{X}_{i}\beta_{2}-\widehat{\mathbf{T}u_{i}}^{S_{b}}\rho)^{\prime}V_{i}^{-1}(\mathbb{T}_{i}y_{i}-\mathbb{T}_{i}x_{1i}\beta_{1}-\mathbb{T}_{i}\tilde{X}_{i}\beta_{2}-\widehat{\mathbf{T}u_{i}}^{S_{b}}\rho)$. Importantly, we can just perform ordinary least squares estimation if $\mathbb{T}_{i}=D_{i}$, $T=2$, and $\omega_{it}$ is homoskedastic and uncorrelated across both $i$ and $t$.
@@
    Estimate the conditional expectation $\mathbb{E}[\mathbf{T}x_{1it}|I_{t}]$ using a super learner and data correspondig to $S_{b}^{c}$.
    Predict $\mathbb{E}[\mathbf{T}x_{1it}|I_{t}]$ for $i\in S_{b}$.
    Obtain the residuals $\mathbf{T}u_{it}$ for $i\in S_{b}$ using the prediction of $\mathbb{E}[\mathbf{T}x_{1it}|I_{t}]$ of previous step.
    Estimate $\theta_{o}=[\beta_{1o},\beta_{2o},\rho_{o}]^{\prime}$ solving the sample moment conditions 26 for $i\in S_{b}$.
@@
First-difference transformation (continued):
@@
$\Delta y_{it}=\Delta x_{1it}\beta_{1_{o}}+\Delta\tilde{x}_{it}^{\prime}\beta_{2_{o}}+\rho_{o}\Delta u_{it}+\Delta\omega_{it},i\in\{1,...,N\},t\in\{2,...,T_{i}\}.$ (27)
First, we partition the set $\{1,2,...,N\}$ in B subsets $S_{1}$, $S_{2}$, …, $S_{B}$, and we estimate $\mathbb{E}[\Delta x_{1it}|\tilde{x}_{it},\tilde{x}_{it-1},z_{it},z_{it-1}]$ using a super learner with partition $S_{b}^{c}$. Then, we get the residuals $\widehat{\Delta u}_{it}^{S_{b}}=\Delta x_{1it}-\widehat{\Delta g_{o}^{S_{b}^{c}}(\tilde{x}_{it},z_{it})}$ for partition $S_{b}$. We obtain the estimator of $\theta_{o}$ for partition $S_{b}$ as the solution of the sample moment conditions such that:
$\frac{1}{n_{T,b}}\sum_{i\in S_{b}}\phi(O_{i};\hat{\theta}_{b},\widehat{\Delta g}_{o}^{S_{b}^{c}})=0,$ (28)
where $\phi(O_{i};\hat{\theta}_{b},\widehat{\Delta g}_{o}^{S_{b}^{c}})=(\tilde{H}_{i}^{S_{b}})^{\prime}V_{i}^{-1}(D_{i}y_{i}-\tilde{H}_{i}^{S_{b}}\hat{\theta}_{b}))$, $\tilde{H}_{i}^{S_{b}}=[D_{i}x_{1i}\quad D_{i}\tilde{x}_{i}\quad\widehat{\Delta u}_{i}^{S_{b}}]$, and $\widehat{\Delta u}_{i}^{S_{b}}$ are the residuals obtained in the first stage for partition $S_{b}$.
@@
    Estimate the conditional expectation $\mathbb{E}[\Delta x_{1it}|\tilde{x}_{it},\tilde{x}_{it-1},z_{it},z_{it-1}]$ using a super learner and data correspondig to $S_{b}^{c}$.
    Predict $\mathbb{E}[\Delta x_{1it}|\tilde{x}_{it},\tilde{x}_{it-1},z_{it},z_{it-1}]$ for $i\in S_{b}$.
    Obtain the residuals $u_{it}$ for $i\in S_{b}$ using the prediction of $\mathbb{E}[\Delta x_{1it}|\tilde{x}_{it},\tilde{x}_{it-1},z_{it},z_{it-1}]$ of previous step.
@@
7 Large sample properties of the SLFC estimator
Theorem 1.
Consistency of the SLCF estimator
If $\{O_{i}=(y_{i},X_{i},z_{i})\}_{i}\in[N]$ satisfy assumptions 1 to 4, assumptions of subsection 11.3 in the Annex hold, i) $\Sigma_{i}$ is positive semi-definite matrix and $\Sigma_{i}\mathbb{E}[\phi(O_{i};\theta,\mathbf{T}g_{o})]=0$ only if $\theta=\theta_{o}$, ii) $\theta\in\Theta$ with $\Theta$ compact, iii) $\phi(O_{i};\theta,\mathbf{T}g_{o})$ is continuous at each $\theta\in\Theta$, iv) the function class =$\mathcal{F}=\{\phi(O_{i};\theta,\mathbf{T}g_{o}\}$ satisfies a uniform integrability condition $\mathbb{E}[\operatorname*{sup_{\theta\in\Theta}}\left\|\phi(O_{i};\theta,\mathbf{T}g_{o})\right\|]<\infty$, then $\hat{\theta}_{o}$ is consistent.
The proof is presented in the Annex 11.
@@ 8.2 The results
In this section, we present the results of the simulation experiment. In figure 2, we present the average estimated parameter of interest $\beta_{1o}$ for different values of $a$. The methods are Within OLS (WOLS), Within 2SLS (W2SLS) using polynomial of degree 1 of the covariate and the instrumental variables, Within 2SLS using polynomial of degree 5 of the covariate and the instrumental variables (W2SLS_poly), the Super Learner Control Function estimator on the First-Differenced data (FDSLCF), and the Super Learner Control Function estimator on the Within-transformed data (WSLCF). In addition, we plot the average correlation between the endogenous covariate and the instrumental variable. We also plot the p-value of the estimated control functions using the first-differenced and the within transformed data. As $a$ is larger, the nonlinearity in the reduced form equation is more important. This translates in a lower correlation between the endogenous regressor and the instrumental variables. As a result, the Within-2SLS estimators present a decay in their performance. In contrast, the Super Learner Control Function estimators remain stable around the true value of the parameter of interest $\beta_{1o}=1$.
@@
As an illustration, we estimate the causal effect of air pollution on student performance in the United States. The data on air pollution are obtained from Deryugina et al. (2019). This dataset presents daily data on air pollutants: PM 2.5, ozone ($O_{3}$), carbon monoxide (CO), sulfur dioxide ($SO_{2}$), nitrogen dioxide ($NO_{2}$). It also contains data on wind direction, wind speed, and weather conditions at county level. We aggregate the data to yearly level by averaging across days. The data on student performance are obtained from Stanford Educational Data Archive (SEDA, Fahle et al. 2024) for county-yearly level. We merge both datasets by county-year indicator, and keep subject mathematics and cohort 2006. Our final sample is composed of years 2009 to 2014 and 3065 counties.
We specify the following linear regression for county $d$, at year $t$:
@@ 10 Conclusions
This paper proposes a triangular simultaneous equation model for panel data with additive separable individual-specific fixed effects composed of a linear structural equation with a nonlinear reduced form equation to deal with the issue of weak instrumental variables due to nonlinearities in panel data settings. The parameter of interest is the structural parameter of the endogenous variable. The identification of this parameter is obtained under the assumption of available exclusion restrictions and using a control function approach. Estimating the parameter of interest is done using an estimator that we call Super Learner Control Function estimator (SLCFE). The estimation procedure is composed of two main steps and sample splitting. We estimate the control function with a super learner using sample splitting. In the following step, we use the estimated control function to control for endogeneity in the structural equation. Sample splitting is done across the individual dimension. We perform a Monte Carlo simulation to test the performance of the estimators proposed. We conclude that the estimators perform well, provided that we can learn accurately the nuisance parameter in the first stage.
11 Annex
11.1 Notation
A set $\{1,2,...,A\}$ is denoted as $[A]$. ||$\cdot$|| ⋅ | refers to the cardinality of a set. $\left\|\cdot\right\|_{L^{p}}$ is the $L^{p}$ norm defined on the space $L^{p}(\Omega,P)$. $\left\|\cdot\right\|_{p}$ is the p-norm.
11.2 Definitions
We follow Emmenegger and Bühlmann (2023) and provide the following definitions. Each individual $i$ presents $T_{i}$ observations. We assume that $T_{i}$ is uniformly bounded such that $T_{i}<T_{max}$. The total number of observations is denoted by $N_{T}=\sum_{i}T_{i}$. The number of sample splits $B$ is not random and independent of $N$. The data is defined as $\{O_{i}=(X_{i},z_{i},y_{i})\}_{i\in[N]}$ with $X_{i}=[x_{1i},\tilde{X}_{i}]$, and denote the probability distribution $P$ of the grouped data $O_{i}$. $\{\mathcal{P}_{N}\}_{N\geq 1}$ is a sequence of probability distributions. A partition of the set $[N]$ is $S_{1}$, $S_{2}$, …, $S_{B}$. The total number of observations belonging to a partition $S_{b}\in[N]$ is equal to $n_{T,b}=\sum_{i\in S_{b}}T_{i}$. We assume that the subsets $S_{b}$ have similar size such that $Bn_{T,b}=N_{T}+o(1),\quad\forall b\in[B]$. We denote $S_{b}^{c}=\{O_{i}=(y_{i},X_{i},z_{i})\}_{i\in S_{b}^{c}}$, and the estimator of the nuisance parameter $\widehat{\mathbf{T}g}_{o}^{S_{b}^{c}}$ with data from $S_{b}^{c}$. The empirical mean $\mathbb{E}_{n_{T,b}}=\frac{1}{n_{T_{b}}}\sum_{i\in S_{b}}\phi(O_{i};\theta,\mathbf{T}g)$. We define $\{\delta_{N}\}_{N\geq B}$ as a sequence of positive numbers that converge to 0 as $N\rightarrow\infty$, and $\{\delta_{N}\}\geq N^{-1/2}$.
11.2.1 Population Loss Function
We define the population following population loss function:
$Q(\theta,\mathbf{T}g_{o})=\mathbb{E}[(\mathbb{T}_{i}(y_{i}-H_{i}\theta))^{\prime}V_{i}^{-1}\mathbb{T}_{i}(y_{i}-H_{i}\theta)],$
where $V_{i}=\mathbb{E}[(\mathbb{T}_{i}\omega_{i})(\mathbb{T}_{i}\omega_{i})^{\prime}]$, and $\theta=[\beta_{1},\quad\beta_{2}^{\prime},\quad\rho]^{\prime}$. $V_{i}$ is a scalar if $T_{i}=2,\forall i\in[N]$, $\mathbb{T}_{i}=D_{i}$, and the error term $\omega_{it}$ is homoskedatisc and uncorrelated across $i$ and $t$.
11.2.2 Empirical Loss Function
We define the population following loss function:
$L_{N}(\theta,\mathbf{T}g_{o})=\sum_{i}^{N}\frac{1}{N}[(\mathbb{T}_{i}y_{i}-\mathbb{T}_{i}H_{i}\theta)^{\prime}V_{i}^{-1}(\mathbb{T}_{i}y_{i}-\mathbb{T}_{i}H_{i}\theta)].$
11.2.3 Population moment condition
We define the population moment condition:
$m(O_{i};\theta,\mathbf{T}g_{o})=\mathbb{E}[(\mathbb{T}_{i}y_{i}-\mathbb{T}_{i}H_{i}\theta)^{\prime}V_{i}^{-1}\mathbb{T}_{i}H_{i}]$
with $m(O_{i};\theta_{o},\mathbf{T}g_{o})=0$.
11.2.4 Score function
The score function is $\phi(O_{i};\theta,\mathbf{T}g_{o})=(\mathbb{T}_{i}y_{i}-\mathbb{T}_{i}H_{i}\theta)^{\prime}V_{i}^{-1}\mathbb{T}_{i}H_{i}$.
11.2.5 Sample moment condition
We define the sample moment condition:
$\hat{m}(O_{i};\theta,\mathbf{T}g_{o})=\frac{1}{N}\sum_{i}(\mathbb{T}_{i}y_{i}-\mathbb{T}_{i}H_{i}\theta)^{\prime}V_{i}^{-1}\mathbb{T}_{i}H_{i}.$
11.3 Assumptions
Assumption 5.
$\theta\in\Theta$, the set $\Theta$ is bounded, contains $\theta_{o}$.
This assumption is similar to assumption 4.1 of Emmenegger and Bühlmann (2023).
Assumption 6.
The set $\mathcal{G}$ consists P-integrable functions $\mathbf{T}g$ with pth moment that exists and it contains $\mathbf{T}g_{o}$, and
$\left\|\mathbf{T}g-\mathbf{T}g_{o}\right\|_{P,2}\leq\delta_{N}.$
Assumption 7.
For all N, all $i\in[N]$, all $P\in\mathcal{P}_{N}$, all $b\in[B]$, and $p>8$ we have:
  1. 7.1
    At the true $\theta_{0}$ and the true $g_{0}$, the data $\{O_{i}=(W_{i},X_{i},Z_{i},Y_{i})\}_{i\in[N]}$ satisfies the identifiability condition.
    $\mathbb{E}[\mathbb{E}_{n_{T,b}}\phi(O_{i};\theta_{o},\mathbf{T}g_{o})]=0.$
  2. 7.2
    $\mathbb{E}[\mathbb{E}_{n_{T,b}}\phi(O_{i};\theta_{o},\mathbf{T}g_{o})\phi(O_{i};\theta_{o},\mathbf{T}g_{o})^{\prime}]=\Sigma_{i}$ with $\Sigma_{i}$ a semi positive-definite matrix, with eigenvalues bounded away from 0 and finite, with the minimum eigenvalue $\lambda_{T}>0$, and the maximum eigenvalue $\lambda_{1}<\infty$.
  3. 7.3
    There exists finite real constants $C_{1}$,$C_{2}$, $C_{3}$, $C_{4}$ satisfying $||y_{i}||_{P,p}\leq C_{1}$, $||z_{i}||_{P,p}\leq C_{2}$, $||x_{1,i}||_{P,p}\leq C_{3}$, $||\tilde{x}_{i}||_{P,p}\leq C_{4}$ for all $k>2$.
Assumption 7.1 states that the true parameter vector is identifiable. Assumption 7.3 state that the data has finite moments.
This assumption is equivalent to Assumption 5.1 of Emmenegger and Bühlmann (2023), and Newey and McFadden (1994).
11.4 Consistency
The proof of consistency of $\hat{\theta}_{o}$ follows Emmenegger and Bühlmann (2023), and Newey and McFadden (1994).
Proof 2.
Proof of Theorem 1
Condition i) is satisfied by Assumptions 7.1, and 7.2, condition ii) is satisfied by Assumption 5, condition iii) and iv) are satisfied by lemma 12, lemma 15 of Emmenegger and Bühlmann (2023), and lemma 1. This implies that the estimator $\hat{\theta}$ is consistent (Emmenegger and Bühlmann, 2023).
Lemma 1.
We prove that lemma 13 of Emmenegger and Bühlmann (2023) holds in our set up.
Proof 3.
$\begin{split}\phi(O_{i};\theta,\mathbf{T}g)-\phi(O_{i};\theta,\mathbf{T}g_{o})&=(\mathbb{T}_{i}y_{i}-\tilde{H}_{i}\theta)^{\prime}V_{i}^{-1}\tilde{H}_{i}-(\mathbb{T}_{i}y_{i}-\mathbb{T}_{i}H_{i}\theta)^{\prime}V_{i}^{-1}\mathbb{T}_{i}H_{i}\\ &=(\mathbb{T}_{i}y_{i})^{\prime}V_{i}^{-1}[\tilde{H}_{i}-\mathbb{T}_{i}H_{i}]+[(\tilde{H}_{i}-\mathbb{T}_{i}H_{i})\theta)]^{\prime}V_{i}^{-1}[(\tilde{H}_{i}-\mathbb{T}_{i}H_{i})],\end{split}$ (38)
Now, notice that $(\tilde{H}_{i}-\mathbb{T}_{i}H_{i})=[\mathbb{T}_{i}x_{1i},\mathbb{T}_{i}\tilde{X}_{i},\widehat{\mathbf{T}u}_{i}]-[\mathbb{T}x_{1i},\mathbb{T}\tilde{X}_{i},\mathbb{T}u_{i}]=[0,0,\widehat{\mathbf{T}u}_{i}-\mathbb{T}_{i}u_{i}]=[0,0,\xi_{i}]$. Then, we are left with:
$(\mathbb{T}_{i}y_{i})^{\prime}V_{i}^{-1}\xi_{i}+\theta^{\prime}\xi_{i}^{\prime}V_{i}^{-1}\xi_{i}.$ (39)
Then:
$\begin{split}\operatorname*{sup}_{\theta\in\Theta,\mathbf{T}g\in\mathcal{G}}\left\|\mathbb{E}_{P}[\mathbb{E}_{n_{T,b}}[(\mathbb{T}_{i}y_{i})^{\prime}V_{i}^{-1}\xi_{i}+\theta^{\prime}\xi_{i}^{\prime}V_{i}^{-1}\xi_{i}]]\right\|&\leq\operatorname*{sup}_{\theta\in\Theta,\mathbf{T}g\in\mathcal{G}}\left\|\mathbb{E}_{P}[(\mathbb{T}_{i}y_{i})^{\prime}V_{i}^{-1}\xi_{i}]\right\|+\operatorname*{sup}_{\theta\in\Theta,\mathbf{T}g\in\mathcal{G}}\left\|\mathbb{E}_{P}[\theta^{\prime}\xi_{i}^{\prime}V_{i}^{-1}\xi_{i}]\right\|\\ &\leq\operatorname*{sup}_{\theta\in\Theta,\mathbf{T}g\in\mathcal{G}}\left\|\mathbb{E}_{P}[\sum_{t}\lambda_{t}(q_{t}^{\prime}\mathbb{T}_{i}y_{i})(q_{t}^{\prime}\xi_{i})]\right\|+\\ &\operatorname*{sup}_{\theta\in\Theta,\mathbf{T}g\in\mathcal{G}}\left\|\theta\mathbb{E}_{P}[\sum_{t}\lambda_{t}(q_{t}^{\prime}\xi_{i}^{\prime})(q_{t}^{\prime}\xi_{i})]\right\|\\ &\leq\operatorname*{sup}_{\theta\in\Theta,\mathbf{T}g\in\mathcal{G}}\left\|\mathbb{E}_{P}[\lambda_{1}\sum_{t}(q_{t}^{\prime}\mathbb{T}_{i}y_{i})(q_{t}^{\prime}\xi_{i})]\right\|+\\ &\operatorname*{sup}_{\theta\in\Theta,\mathbf{T}g\in\mathcal{G}}[\lambda_{1}\left\|\theta\mathbb{E}_{P}\sum_{t}(q_{t}^{\prime}\xi_{i})(q_{t}^{\prime}\xi_{i})]\right\|\\ &\leq\operatorname*{sup}_{\theta\in\Theta,\mathbf{T}g\in\mathcal{G}}\left\|\lambda_{1}\left\|\mathbb{T}_{i}y_{i}\right\|_{P,2}\left\|\xi_{i}\right\|_{P,2}\right\|+\\ &\operatorname*{sup}_{\theta\in\Theta,\mathbf{T}g\in\mathcal{G}}\lambda_{1}\left\|\theta\right\|\left\|\xi_{i}\right\|_{P,2}\end{split}$ (40)
Since $\Theta$ is bounded, $\left\|\mathbf{T}g-\mathbf{T}g_{o}\right\|_{P,2}<\delta_{N}$, the last expression is bounded by $\delta_{N}$ by assumptions 5, and 7.